# Solving Fredholm Integral Equations of the Second Kind via Wasserstein Gradient Flows

### v2

@@ E.2 Proof of Proposition 6
First, we recall that $b$ is given for any $x\in\mathbb{R}^{d}$, $\nu\in\mathcal{P}(\mathbb{R}^{d})$ and $y\in\mathbb{R}^{p}$ by
$b(x,\nu)=\int b^{\eta}(\bm{X}_{t},z,\pi_{t})\mathrm{d}\pi_{t}(z)-\alpha\nabla U(\bm{X}_{t})$
@@
Furthermore, using Assumption 1 we have for any $x_{1},x_{2}\in\mathbb{R}^{d}$, $\pi_{1},\pi_{2}\in\mathcal{P}(\mathbb{R}^{d})$ and $y\in\mathbb{R}^{p}$,
@@
$\leq\lambda\mathtt{M}/\eta+\lambda(\mathtt{M}+1)/\eta=:C_{0}.$
@@ E.3 Proof of Proposition 7
$+(1+\alpha)\int_{s}^{t}\left\|\bm{B}_{u}^{1}\right\|^{2}\mathrm{d}u,$
from which it follows that
@@
$-(1+\alpha)dt+(1+\alpha)\int_{s}^{t}\left\|\bm{B}_{u}^{1}\right\|^{2}\mathrm{d}u.$
Denote by $(\mathcal{F}_{t}^{N})_{t\geq 0}$ the natural filtration of $\{(\bm{B}_{t}^{k})_{t\geq 0}\}_{k=1}^{N}$. Then,
$\mathbb{E}\left[\bm{M}_{t}^{N}\mid\mathcal{F}_{s}^{N}\right]$ $=\frac{1}{2}\left\|\bm{X}_{s}^{1,N}\right\|^{2}-\int\limits_{0}^{s}\langle\bm{X}_{u}^{1,N},b(\bm{X}_{u}^{1,N},\pi^{N}_{u})\rangle\mathrm{d}u$
$-(1+\alpha)dt+(1+\alpha)d(t-s)$
$=\bm{M}_{s}^{N},$
showing that $\bm{M}_{t}^{N}$ is an $\mathcal{F}^{N}$-local martingale.
@@
$\qquad\phantom{=}+\mathbb{E}\left[\int\limits_{s}^{t}\left\{-(1+\alpha)\mathtt{m}\left\|\bm{X}_{u}^{1,N}\right\|^{2}+(1+\alpha)\left\|\nabla U(0)\right\|\left\|\bm{X}_{u}^{1,N}\right\|+(1+\alpha)(\mathtt{c}+d)\right\}\mathrm{d}u\right]$
@@
$\qquad\phantom{=}+\int\limits_{s}^{t}\left\{-(1+\alpha)\mathtt{m}\mathbb{E}\left[\left\|\bm{X}_{u}^{1,N}\right\|^{2}\right]+(1+\alpha)\left\|\nabla U(0)\right\|\mathbb{E}\left[\left\|\bm{X}_{u}^{1,N}\right\|\right]+(1+\alpha)(\mathtt{c}+d)\right\}\mathrm{d}u,$
@@
$-(1+\alpha)\mathtt{m}\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]+(1+\alpha)\left\|\nabla U(0)\right\|\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|\right]$
@@
$\frac{1}{2}\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $C_{0}(\mathcal{V}_{t}^{N})^{1/2}-(1+\alpha)\mathtt{m}\mathcal{V}_{t}^{N}+(1+\alpha)\left\|\nabla U(0)\right\|(\mathcal{V}_{t}^{N})^{1/2}+(1+\alpha)(\mathtt{c}+d)$
$=$ $(C_{0}+(1+\alpha)\left\|\nabla U(0)\right\|)(\mathcal{V}_{t}^{N})^{1/2}+(1+\alpha)(\mathtt{c}+d)-(1+\alpha)\mathtt{m}\mathcal{V}_{t}^{N}$
Noting that for any $\mathsf{a},\mathsf{b}$ such that $\mathsf{ab}\geq 1/2$, for any $x\geq 0$ we have $\sqrt{x}\leq\mathsf{a}+\mathsf{b}x$, and setting $\mathsf{a}=(C_{0}+(1+\alpha)\left\|\nabla U(0)\right\|)/((1+\alpha)\mathtt{m})$ and $\mathsf{b}=1/2\mathsf{a}$ we have:
$\frac{1}{2}\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $\frac{(C_{0}+(1+\alpha)\left\|\nabla U(0)\right\|)^{2}}{(1+\alpha)\mathtt{m}}+(1+\alpha)(\mathtt{c}+d)-\frac{1}{2}(1+\alpha)\mathtt{m}\mathcal{V}_{t}^{N}$
@@
$C=\frac{2}{(1+\alpha)\mathtt{m}}\left[\frac{(C_{0}+(1+\alpha)\left\|\nabla U(0)\right\|)^{2}}{(1+\alpha)\mathtt{m}}+(1+\alpha)(\mathtt{c}+d)\right].$ (54)
Therefore, letting $t\to+\infty$ we get that for any $N\in\mathbb{N}$, $\int\limits_{\mathbb{R}^{d}}\left\|x\right\|^{2}\mathrm{d}\pi^{N}(x)\leq C$. Hence, $\{\pi^{N}\}_{N\in\mathbb{N}}$ is relatively compact in $\mathcal{P}_{1}(\mathbb{R}^{d})$ using [4, Proposition 7.1.5].
Let $\pi^{\star}$ be a cluster point of $\{\pi^{N}\}_{N\in\mathbb{N}}$. Le us denote by $\pi_{t}(\pi^{\star}),\pi_{t}(\pi^{\star})^{N}$ the law of $\bm{X}_{t}$ following (9) with initial condition $\bm{X}_{0}\sim\pi^{\star}$ and $\bm{X}_{t}^{1,N}$ following (11) with initial condition $\bm{X}_{0}^{1,N}\sim\pi^{\star}$. Let $(N_{k})_{k\in\mathbb{N}}$ be an increasing sequence such that $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. We have that for any $t\geq 0$
@@
$+\mathbf{W}_{1}(\pi_{t}(\pi^{N_{k}}),\pi_{t}(\pi^{\star}))+\mathbf{W}_{1}(\pi_{t}(\pi^{\star}),\pi_{\alpha,\eta}^{\star}).$
We now control each of these terms. Let $\varepsilon>0$, for sufficiently large $t\geq 0$ Proposition 4 ensures $\mathbf{W}_{1}(\pi_{t}(\pi^{\star}),\pi_{\alpha,\eta}^{\star})\leq\varepsilon$. Using the Lipschitz continuity of the drift established in (41) and Assumption 3 with [16, Lemma 20], there exists $k_{0}\in\mathbb{N}$ such that for any $k\geq k_{0}$ and any $t\in[0,T]$ we have
@@
since we assumed $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. Using Proposition 5, there exists $k_{1}\in\mathbb{N}$ such that for any $k\geq k_{1}$ we have that $\mathbf{W}_{1}(\pi_{t}^{N_{k}}(\pi^{N_{k}}),\pi_{t}(\pi^{N_{k}}))\leq\varepsilon$. Since $\pi^{N_{k}}$ is invariant for $(\bm{X}_{t}^{1:N_{k}})_{t\geq 0}$ we get that $\mathbf{W}_{1}(\pi^{N_{k}},\pi_{t}^{N_{k}}(\pi^{N_{k}}))=0$. Finally, there exists $k_{2}\in\mathbb{N}$ such that for any $k\geq k_{2}$ we have $\mathbf{W}_{1}(\pi^{\star},\pi^{N_{k_{2}}})\leq\varepsilon$ since we assumed $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. Combining these results in (55), we get that $\mathbf{W}_{1}(\pi^{\star},\pi_{\alpha,\eta}^{\star})\leq 4\varepsilon$. Therefore, since $\varepsilon>0$ is arbitrary, we have that $\pi^{\star}=\pi_{\alpha,\eta}^{\star}$, which concludes the proof.

### v3

@@ E.2 Proof of Proposition 6
First, we recall that $b$ is given for any $x\in\mathbb{R}^{d}$, $\nu\in\mathcal{P}(\mathbb{R}^{d})$ by
$b(x,\nu)=\int b^{\eta}(x,z,\nu)\mathrm{d}\nu(z)-\alpha\nabla U(x)$
@@
Furthermore, using Assumption 1 we have for any $x_{1},x_{2}\in\mathbb{R}^{d}$, $\pi_{1},\pi_{2}\in\mathcal{P}(\mathbb{R}^{d})$,
@@
$\leq\lambda\mathtt{M}/\eta+(\lambda+1)\mathtt{M}/\eta=:C_{0}.$
@@ E.3 Proof of Proposition 7
$+(1+\alpha)d(t-s)+\sqrt{2(1+\alpha)}\int_{s}^{t}\langle\bm{X}_{u}^{1,N},\mathrm{d}\bm{B}_{u}^{1}\rangle,$
so that
@@
$-(1+\alpha)ds+\sqrt{2(1+\alpha)}\int_{s}^{t}\langle\bm{X}_{u}^{1,N},\mathrm{d}\bm{B}_{u}^{1}\rangle$
$=\bm{M}_{s}^{N}+\sqrt{2(1+\alpha)}\int_{s}^{t}\langle\bm{X}_{u}^{1,N},\mathrm{d}\bm{B}_{u}^{1}\rangle$
It follows that
$\mathbb{E}\left[\bm{M}_{t}^{N}-\bm{M}_{s}^{N}\mid\mathcal{F}_{s}^{N}\right]=\sqrt{2(1+\alpha)}\mathbb{E}\left[\int_{s}^{t}\langle\bm{X}_{u}^{1,N},\mathrm{d}\bm{B}_{u}^{1}\rangle\mid\mathcal{F}_{s}^{N}\right].$
[…]
$\left\|b(x,\nu)\right\|\leq C(1+\left\|x\right\|).$
Observe that
$\left\|b(x,\nu)\right\|=\left\|b(x,\nu)-b(0,\nu)+b(0,\nu)\right\|\leq C_{1}\left\|x\right\|+\left\|b(0,\nu)\right\|$
due to Lipschitzness. Moreover, $\left\|b(0,\nu)\right\|$ is finite :
$\left\|b(0,\nu)\right\|$ $\leq\int\left\|b^{\eta}(0,z,\nu)\right\|\mathrm{d}\nu(z)+\alpha\left\|\nabla U(0)\right\|$
$\leq C_{0}+\alpha\left\|\nabla U(0)\right\|$
[…]
@@
$\qquad\phantom{=}+\mathbb{E}\left[\int\limits_{s}^{t}\left\{-\alpha\mathtt{m}\left\|\bm{X}_{u}^{1,N}\right\|^{2}+\alpha\left\|\nabla U(0)\right\|\left\|\bm{X}_{u}^{1,N}\right\|+(1+\alpha)(\mathtt{c}+d)\right\}\mathrm{d}u\right]$
@@
$\qquad\phantom{=}+\int\limits_{s}^{t}\left\{-\alpha\mathtt{m}\mathbb{E}\left[\left\|\bm{X}_{u}^{1,N}\right\|^{2}\right]+\alpha\left\|\nabla U(0)\right\|\mathbb{E}\left[\left\|\bm{X}_{u}^{1,N}\right\|\right]+(1+\alpha)(\mathtt{c}+d)\right\}\mathrm{d}u,$
@@
$-\alpha\mathtt{m}\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|^{2}\right]+\alpha\left\|\nabla U(0)\right\|\mathbb{E}\left[\left\|\bm{X}_{t}^{1,N}\right\|\right]$
@@
$\frac{1}{2}\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $C_{0}(\mathcal{V}_{t}^{N})^{1/2}-\alpha\mathtt{m}\mathcal{V}_{t}^{N}+\alpha\left\|\nabla U(0)\right\|(\mathcal{V}_{t}^{N})^{1/2}+(1+\alpha)(\mathtt{c}+d)$
$=$ $(C_{0}+\alpha\left\|\nabla U(0)\right\|)(\mathcal{V}_{t}^{N})^{1/2}+(1+\alpha)(\mathtt{c}+d)-\alpha\mathtt{m}\mathcal{V}_{t}^{N}$
Noting that for any $\mathsf{a},\mathsf{b}$ such that $\mathsf{ab}\geq 1/2$, for any $x\geq 0$ we have $\sqrt{x}\leq\mathsf{a}+\mathsf{b}x$, and setting $\mathsf{a}=(C_{0}+\alpha\left\|\nabla U(0)\right\|)/(\alpha\mathtt{m})$ and $\mathsf{b}=1/2\mathsf{a}$ we have:
$\frac{1}{2}\mathrm{d}\mathcal{V}_{t}^{N}/\mathrm{d}t\leq$ $\frac{(C_{0}+\alpha\left\|\nabla U(0)\right\|)}{\alpha\mathtt{m}}+(1+\alpha)(\mathtt{c}+d)-\frac{1}{2}\alpha\mathtt{m}\mathcal{V}_{t}^{N}$
@@
$C=\frac{2}{\alpha\mathtt{m}}\left[\frac{(C_{0}+\alpha\left\|\nabla U(0)\right\|)}{\alpha\mathtt{m}}+(1+\alpha)(\mathtt{c}+d)\right].$ (53)
Therefore, letting $t\to+\infty$ we get that for any $N\in\mathbb{N}$, $\int\limits_{\mathbb{R}^{d}}\left\|x\right\|^{2}\mathrm{d}\pi^{N}(x)\leq C$. Hence, $\{\pi^{N}\}_{N\in\mathbb{N}}$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R}^{d})$ using [4, Proposition 7.1.5].
Let $\pi^{\star}$ be a cluster point of $\{\pi^{N}\}_{N\in\mathbb{N}}$. Le us denote by $\pi_{t}(\pi^{\star}),\pi_{t}^{N}(\pi^{\star})$ the law of $\bm{X}_{t}$ following (9) with initial condition $\bm{X}_{0}\sim\pi^{\star}$ and $\bm{X}_{t}^{1,N}$ following (11) with initial condition $\bm{X}_{0}^{1,N}\sim\pi^{\star}$. Let $(N_{k})_{k\in\mathbb{N}}$ be an increasing sequence such that $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. We have that for any $t\geq 0$
@@
$+\mathbf{W}_{1}(\pi_{t}(\pi^{N_{k}}),\pi_{t}(\pi^{\star}))+\mathbf{W}_{1}(\pi_{t}(\pi^{\star}),\pi_{\alpha,\eta}^{\star}),$
where $\pi^{N_{k}}$ denotes the invariant measure of $(\bm{X}_{t}^{1:N_{k}})_{t\geq 0}$ which exists by Proposition 6. We now control each of these terms. Let $\varepsilon>0$, for sufficiently large $t\geq 0$ Proposition 4 ensures $\mathbf{W}_{1}(\pi_{t}(\pi^{\star}),\pi_{\alpha,\eta}^{\star})\leq\varepsilon$. Using the Lipschitz continuity of the drift established in (41) and Assumption 3 with [16, Lemma 20], there exists $k_{0}\in\mathbb{N}$ such that for any $k\geq k_{0}$ and any $t\in[0,T]$ we have
@@
since we assumed $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. Using Proposition 5, there exists $k_{1}\in\mathbb{N}$ such that for any $k\geq k_{1}$ we have that $\mathbf{W}_{1}(\pi_{t}^{N_{k}}(\pi^{N_{k}}),\pi_{t}(\pi^{N_{k}}))\leq\varepsilon$. Since $\pi^{N_{k}}$ is invariant for $(\bm{X}_{t}^{1:N_{k}})_{t\geq 0}$ we get that $\mathbf{W}_{1}(\pi^{N_{k}},\pi_{t}^{N_{k}}(\pi^{N_{k}}))=0$. Finally, there exists $k_{2}\in\mathbb{N}$ such that for any $k\geq k_{2}$ we have $\mathbf{W}_{1}(\pi^{\star},\pi^{N_{k}})\leq\varepsilon$ since we assumed $\lim_{k\to+\infty}\mathbf{W}_{1}(\pi^{N_{k}},\pi^{\star})=0$. Combining these results in (54), we get that $\mathbf{W}_{1}(\pi^{\star},\pi_{\alpha,\eta}^{\star})\leq 4\varepsilon$. Therefore, since $\varepsilon>0$ is arbitrary, we have that $\pi^{\star}=\pi_{\alpha,\eta}^{\star}$, which concludes the proof.